OpenStudy (anonymous):

In a US household, the voltage in an electric outlet (in volts) is given by the following equation, where t is in seconds. V = 156sin(120πt) However, in a European house, the voltage is given (in the same units) by the following equation. V = 339sin(100πt) Answer the following questions comparing the voltages in the two regions. (a) How much greater is the maximum voltage in a European house than a US house? (b) What is the number of cycles (oscillations) per second in a US house?

OpenStudy (anonymous):

A.) (Amplitude of the European house) - (Amplitude of an American house) = difference. 339-156 = 183. B.) Since the question asks for oscillations the answer will be a frequency. Frequency = 1/Period. The period of the American house is 2pi/120pi which is equal to 1/60. Frequency =1/Period so 1/1/60 = 60. 60 oscillations per second.

OpenStudy (anonymous):

thanks for your help